# Memory Storyboard: Leveraging Temporal Segmentation for Streaming Self-Supervised Learning from Egocentric Videos

### v1

@@ Abstract
Self-supervised learning holds the promise to learn good representations from real-world continuous uncurated data streams. However, most existing works in visual self-supervised learning focus on static images or artificial data streams. Towards exploring a more realistic learning substrate, we investigate streaming self-supervised learning from long-form real-world egocentric video streams. Inspired by the event segmentation mechanism in human perception and memory, we propose “Memory Storyboard” that groups recent past frames into temporal segments for more effective summarization of the past visual streams for memory replay. To accommodate efficient temporal segmentation, we propose a two-tier memory hierarchy: the recent past is stored in a short-term memory, and the storyboard temporal segments are then transferred to a long-term memory. Experiments on real-world egocentric video datasets including SAYCam [55] and KrishnaCam [52] show that contrastive learning objectives on top of storyboard frames result in semantically meaningful representations which outperform those produced by state-of-the-art unsupervised continual learning methods.
Figure 1: Our proposed Memory Storyboard framework for streaming SSL from egocentric videos. Similar frames are being clustered into temporal segments and their labels (text information for illustration purpose only) are updated in the long-term memory buffer for replay. SSL involves contrastive learning at both the frame and temporal segment levels.
@@ 1 Introduction
Humans are capable of learning continuously from a stream of unlabeled and uncurated perceptual inputs, such as video data, without needing to iterate through multiple exposures or epochs. Since early infancy, humans have accumulated knowledge about the world through a continuous flow of raw visual observations. This capability contrasts sharply with the training paradigm of current methods in self-supervised learning (SSL). While recent SSL approaches have made great strides in learning from large unlabeled datasets [11, 19, 12, 9, 7, 25, 3, 26], they still predominantly rely on static and curated image datasets, such as ImageNet [13], and needs multiple epochs of training for effective learning. This difference in paradigm raises a compelling question: how can we learn good visual representations in a streaming setting—learning from visual inputs in their original temporal order without cycling back?
@@
Motivated by the differences of mechanisms between human learning and standard SSL, we aim to build learning algorithms that can efficiently learn visual representations and concepts from streaming video. One especially relevant mechanism in the human brain is event segmentation [37, 70], where we spontaneously segment visual streams into hierarchically structured events and identify the event boundaries. Take your recent vacation trip as an example—you probably remember separate events and activities like exploring a city, dining at a local restaurant, or relaxing at the beach. The event segmentation mechanism helps us organize memories, recall specific moments, and summarize from lengthened experiences [68, 69].
@@
Drawing inspiration from the way we organize our memory in the brain, we introduce Memory Storyboard, a novel approach for streaming self-supervised learning. Memory Storyboard features a temporal segmentation module, which groups video frames into semantically meaningful temporal segments, resembling the automatic event segmentation of human cognition. Through our temporal contrastive learning objective, these temporal segments effectively facilitate representation learning in streaming videos. To accommodate efficient temporal segmentation, we propose a two-tier hierarchical memory: temporal segmentation in the short-term memory are used to update the temporal class labels in the long-term memory, and a training batch consists of samples mixed from both memories. A high-level diagram of the algorithm is shown in Figure 1.
@@
We conduct experiments on the SAYCam [55] and KrishnaCam [52] datasets of real-world egocentric videos. Memory Storyboard outperforms state-of-the-art unsupervised continual learning methods on downstream image classification and object detection tasks, and significantly reduces the gap between streaming learning and the less flexible IID learning that requires persistent storage of the entire prior video data. We also experiment with different buffer sizes and batch sizes and offer insights on the optimal training batch composition under different memory constraints.
@@
    We demonstrate that Memory Storyboard achieves state-of-the-art downstream performance when trained on real-world egocentric video datasets. Among all the streaming self-supervised learning methods we evaluated, Memory Storyboard is the only one that is competitive with or even outperforms IID training when trained on these datasets.
@@
    We study the effects of training factors including subsampling rate, average segment length, memory buffer size and training batch composition. These studies provide insight for more efficient streaming learning from videos. In particular, we explore the optimal composition ratio of the training batch from short-term vs. long-term memory, under different memory constraints. Larger batches from long-term memory improve performance when we can afford a large memory bank, while smaller batches can help prevent overfitting when we have a small memory bank.
@@ 2 Related Work
Unsupervised Continual Learning (UCL) [44, 53] aims at learning a good representation through an unlabeled non-stationary data stream. Existing works in UCL, notably CaSSLe [17] and Osiris [73], assume that the data stream is composed of a series of episodes, and a stationary data distribution within each episode. This is not as naturalistic and human-like as our streaming setting, where the data distribution changes continuously in through the data stream, and each image appear in the data stream only once. Meanwhile, we showed that existing UCL methods are also effective in our streaming video setting, and can be used together with the supervised contrastive objective.
Self-Supervised Learning.
A large number of self-supervised representation learning methods in computer vision follows the contrastive learning framework [39, 36, 56, 26, 11, 12] which maximizes the agreement of representations of two augmented views of the same image and minimizes that of different images. Extending this idea, the supervised contrastive (SupCon) method [30] uses the labels as an extra supervision signal to get multiple positive crops for each anchor image. Other recent self-supervised learning works include pretext tasks [14, 38, 18, 41], feature space clustering [8, 9, 45], distillation with asymmetric architectures [19, 12], redundancy reduction [71, 7], and masked autoencoding [25]. Most relevant of these to our work, Orhan et al., [40] proposes the temporal classification objective, which outperforms contrastive learning objectives on the SAYCam dataset [55]. Our work enhances the temporal classification method with using a more flexible supervised contrastive objective, and leveraging temporal segmentation [42, 1], which have been used extensively in video summarization [74, 72, 48].
@@
While a number of recent papers have studied streaming learning from images [22, 24, 23, 6], limited works have investigated the problem of streaming learning from a continuous video stream. Roady et al., [47] introduces a benchmark for streaming classification and novelty detection from videos. Zhuang et al., [75] benchmarks many self-supervised learning methods in real-time and life-long learning settings in streaming video, assuming infinite replay buffer size which is unrealistic. Most similar to our setup, Purushwalkam et al., [43] studies the task of continuous representation learning with a SimSiam objective [12] and proposes using a minimum-redundancy replay buffer. Their work also belongs to the broader range of works that study replay buffer sampling strategies in continual learning [2, 62, 57, 21]. Our work extends these prior works by adopting a two-tier replay buffer and a temporal segmentation component. Also relevant to our work, Carreira et al., [10] studies online learning from a continuous video stream with pixel-to-pixel modeling, but their exploration mainly focuses on settings without data augmentation and replay, limiting the efficacy of their framework.
Temporal Segmentation in Human Cognition.
Prior research in psychology and cognitive sciences has shown that humans, including infants, are able to identify boundaries between action segments [37, 70, 5, 50, 4, 65]. Evidence in neuro-imaging further show that event segmentation is an automatic component in human perception [69]. Temporal event segmentation has proven to be critical for memory formation and retrieval [32, 16, 15, 51, 49]. The temporal segmentation component in our proposed framework is motivated by how humans interpret videos as segments with coherent semantics. We demonstrate that temporal segmentation can improve the learned visual representation.
3 Streaming SSL from Egocentric Videos
In streaming self-supervised learning, the goal is to learn useful visual representations from a continuous stream of inputs $(x_{1},x_{2},\dots)$. Here, we focus on the setting where inputs are uniformly sampled from a video stream. Similar to continual learning, we impose a memory budget so that storing the entire video would violate the constraint. Different from standard continual learning, there is no explicit notion of task, and the data distribution shift follows directly from the scene transitions of a video. The learner needs to make changes to the model as it sees new inputs, and finishes learning as soon as it receives the last input of the stream. The streaming setting is similar to Online Continual Learning [33, 20, 61], but the focus here is primarily on streaming video frames instead of a fixed dataset of static images.
Incorporating Training Batches.
We use $x_{start:end}$ to denote the batch of $(end-start)$ images between $x_{start}$ and $x_{end}$. At each training step $t$, the model fetches a new batch of $b$ images $X_{curr}=x_{tb:(t+1)b}$ from the video stream. The model produces model updates on its parameters $\theta_{t+1}$ upon receiving $X_{curr}$. At the end of the video, we evaluate the final model checkpoint $\theta_{T}$ on various downstream tasks such as object classification and detection, which are fundamental tasks for visual scene understanding as they enable models to recognize and interpret the contents of complex environments.
Standard SSL Fails on Streaming Video.
Directly applying the SSL method on $X_{curr}$ gives very poor performance [43, 45]. This is mainly due to two reasons:
  1. 1)
    the non-stationary distribution of visual features in the stream, similar to the catastrophic forgetting problem [34] in supervised continual learning;
  2. 2)
    the high temporal correlation of images in the stream (illustrated in Figure 1). This temporal correlation breaks the IID assumption held by common optimization algorithms like SGD or Adam [31]. For contrastive learning algorithms like SimCLR [11], the similarity across different frames in the same training batch would violate the assumption that each image is different.
Memory Replay.
Similar to previous works [28, 67, 43], we use a replay buffer $M$ with finite size $|M|$ to mitigate these issues. The model can store some of the fetched images in the replay buffer, and use both samples from the replay buffer and the new frames to form a training batch of size $B$. By sampling from the replay buffer we de-correlate the frames in the training batch and at the same time reduce distribution shift between training batches.
@@ 4 Memory Storyboard
We present Memory Storyboard, an effective method for streaming SSL from egocentric videos. Memory Storyboard includes a temporal segmentation module and a two-tier memory hierarchy. It combines a standard self-supervised contrastive loss with a temporal contrastive objective with leverages the temporal class labels produced by the temporal segmentation module. Figure 2 illustrates the details of our method. The overall data processing and training procedure is summarized in Algorithm 2.
@@
We describe our temporal segmentation algorithm as follows. Similar to Potapov et al., [42], we are given a down-sampled video frame sequence $x_{1},x_{2},\cdots,x_{L}$, and a feature extractor $f_{\theta}$. We aim to find change points $t_{1},t_{2},\cdots,t_{n-1}$ so that the video is divided into $n$ semantically-consistent segments $[x_{1},x_{t_{1}}],[x_{t_{1}},x_{t_{2}}],\cdots,[x_{t_{n-1}},x_{L}]$. We also define $t_{0}=0$ and $t_{n}=L$. In this work, we determine the number of segments with $n=\frac{L}{T}$, where $T$ refers to the average segment length and is a hyper-parameter.
@@
The optimization objective of our segmentation algorithm is to maximize the average within-class similarity, i.e.
$\max_{t_{1},t_{2},\cdots,t_{n-1}}\sum_{i=1}^{n}\frac{1}{t_{i}-t_{i-1}}\sum_{j=t_{i-1}}^{t_{i}}\sum_{k=j}^{t_{i}}sim(x_{j},x_{k}).$ (1)
@@
where $sim(x_{j},x_{k})$ denotes the cosine similarity between the embeddings $f_{\theta}(x_{j})$ and $f_{\theta}(x_{k})$. We compute the approximate solution to this optimization problem with a greedy approach, as detailed in Algorithm 1. We adopt this simple temporal segmentation approach in order to get good segmentation results in the beginning, when encoder network does not provide good representations. We leave it to future work for investigating different temporal segmentation strategies.
@@
loss = TCL_loss(data) + CL_loss(data)
@@
Figure 2: Details of our two-tier memory in Memory Storyboard. Long-term memory is updated with reservoir sampling, and short-term memory with first-in-first-out (FIFO). Temporal segmentation is applied on the short-term memory, which then updates the labels of corresponding images in the long-term memory.
Two-tier Memory Hierarchy.
To accommodate efficient temporal segmentation, we propose a two-tier memory hierarchy. Shown in Figure 2, the system includes a long-term memory $M_{long}$ updated with reservoir sampling [58], and a short-term memory storyboard $M_{short}$ updated with a first-in-first-out (FIFO) strategy. We store the temporal index and the temporal class of each frame along with the image in the memory. The short-term memory size $|M_{short}|$ is much smaller than the long-term memory size $|M_{long}|$, allowing efficient temporal segmentation of the recent past. The change points produced by the temporal segmentation component on $M_{short}$ are then used to update the temporal class labels in $M_{long}$.
To increase the horizon of the memory storyboard, we subsample the frames coming from the current stream before adding it to $M_{short}$. The subsampling also reduces the temporal correlation between the frames in the training batch sampled from $M_{short}$. Compared to using a single replay buffer as memory, the two-tier memory hierarchy helps avoid overfitting on the replay buffer and makes sure that the new frames are seen by the model.
Temporal Contrastive Loss.
To effectively utilize the temporal class labels for representation learning, we adopt the supervised contrastive (SupCon) loss [30], which takes the samples with the same temporal class label in a batch as positives and contrasts them from the remainder of the batch. Let $f_{proj}$ be a projector network. For a batch of images with size $B$, we take two random augmentations of each image to get an augmented batch $\tilde{x_{1}},\tilde{x_{2}},\dots,\tilde{x_{2B}}$, and compute $z_{i}=f_{proj}(f_{\theta}(\tilde{x_{j}}))$ be the projected features of each augmented image $\tilde{x_{i}}$. Let $y_{i}$ be the temporal class label of $\tilde{x_{i}}$ and $P(i)=\{p\in\{1,2,\dots,2B\}\backslash\{i\}:y_{p}=y_{i}\}$.
We refer to this as the temporal contrastive loss. It is conceptually similar to the temporal classification loss proposed in [40]. However, in the temporal classification loss, the size of the classification layer needs to be gradually expanded as more data is processed by the model and more temporal classes are formed. Hence the temporal contrastive loss is more flexible and more suitable for the streaming SSL setting.
Overall Loss Function.
In addition to the temporal contrastive loss, we also incorporate a standard self-supervised contrastive loss $\mathcal{L}_{CL}$. In particular, we experimented with the SimCLR loss [11, 54] and the SimSiam loss [12] because they were shown to work well in lifelong self-supervised learning in prior works [75, 43]. The overall loss function is a sum of the temporal contrastive loss and the self-supervised contrastive loss (Equation 3).
$\mathcal{L}=\mathcal{L}_{TCL}+\mathcal{L}_{CL}.$ (3)
Method mini-ImageNet ImageNet Labeled-S mini-ImageNet ImageNet Labeled-S
SVM kNN Linear SVM kNN SVM kNN Linear SVM kNN
IID SimCLR [11] 44.04 34.72 30.44 59.50 59.08 44.04 34.72 30.44 59.50 59.08
IID SimSiam [12] 29.02 31.60 20.92 42.71 56.84 29.02 31.60 20.92 42.71 56.84
SimCLR No Replay 5.76 6.04 2.22 19.13 21.99 5.76 6.04 2.22 19.13 21.99
SimSiam No Replay 6.44 6.32 1.47 22.03 23.16 6.44 6.32 1.47 22.03 23.16
Replay - 10k Replay - 50k
Osiris [73] 31.16 24.74 19.48 45.81 50.12 36.90 28.84 23.16 50.88 52.19
TC [40] 33.92 24.28 19.03 48.09 54.29 36.68 24.08 22.72 52.22 54.71
SimCLR [11] 33.02 26.82 20.13 49.29 53.57 37.96 30.92 23.75 53.67 56.15
   +MinRed [43] 33.62 27.38 20.21 48.88 52.33 38.66 31.26 24.10 54.75 55.85
[…]
SimSiam [12] 20.90 27.02 13.72 39.12 54.09 26.66 30.04 14.44 43.09 56.95
   +MinRed [43] 22.68 27.12 17.85 39.78 51.84 25.58 27.42 18.99 40.37 53.01
[…]
Table 1: Results on streaming SSL from SAYCam [55]. Downstream evaluation on object classification for SSL models trained under the streaming setting. For “No Replay” and “IID” the results are the same for different memory buffer sizes. The “IID” methods are not under the streaming setting and are for reference only as a performance “upper bound” with the same number of gradient updates. Unless specified, standard reservoir sampling is used in the replay buffer.
Method mini-ImageNet ImageNet OAK mini-ImageNet ImageNet OAK
SVM kNN Linear AP50 AP75 SVM kNN Linear AP50 AP75
IID SimCLR [11] 36.90 27.54 23.77 39.54 21.84 36.90 27.54 23.77 39.54 21.84
IID SimSiam [12] 28.58 25.12 22.28 44.86 31.13 28.58 25.12 22.28 44.86 31.13
SimCLR No Replay 4.84 5.16 1.35 14.01 4.67 4.84 5.16 1.35 14.01 4.67
SimSiam No Replay 8.88 7.52 1.92 27.34 9.47 8.88 7.52 1.92 27.34 9.47
Replay - 10k Replay - 50k
Osiris [73] 30.10 23.94 19.03 32.25 16.20 32.38 24.46 20.85 33.78 17.91
TC [40] 32.58 24.24 19.19 32.61 17.41 32.94 24.84 20.50 28.56 12.29
SimCLR [11] 31.46 24.70 19.09 31.92 17.40 34.98 27.14 22.37 33.30 14.92
   +MinRed [43] 31.56 24.30 19.93 34.78 19.15 34.84 27.16 22.29 35.65 19.94
[…]
SimSiam [12] 19.16 18.84 12.94 39.38 27.31 21.84 20.18 14.13 41.13 28.62
   +MinRed [43] 20.90 18.28 14.53 43.74 27.35 22.88 20.36 17.64 44.17 29.36
[…]
Table 2: Results on streaming SSL from KrishnaCam [52]. Downstream evaluation on object classification and object detection for SSL models trained on under the streaming setting. For “No Replay” and “IID” the results are the same for different memory buffer sizes. The “IID” methods are not under the streaming setting and are for reference only as a performance “upper bound” with the same number of gradient updates. Unless specified, standard reservoir sampling is used in the replay buffer.
Warm-Start Training.
At the beginning of training, the model has only seen a very limited amount of data from the video stream. Even with a memory buffer, there is likely high temporal correlation between the sampled frames and could cause instability in the training. To alleviate this problem, we warm-start the system by making no model updates on the first $M_{long}$ frames of the stream and just use them to fill the memory. The warm-start phase ensures that the model is trained on de-correlated samples from the buffer starting from the beginning.

### v3

@@ Abstract
Self-supervised learning holds the promise of learning good representations from real-world continuous uncurated data streams. However, most existing works in visual self-supervised learning focus on static images or artificial data streams. Towards exploring a more realistic learning substrate, we investigate streaming self-supervised learning from long-form real-world egocentric video streams. Inspired by the event segmentation mechanism in human perception and memory, we propose “Memory Storyboard,” a novel continual self-supervised learning framework that groups recent past frames into temporal segments for a more effective summarization of the past visual streams for memory replay. To accommodate efficient temporal segmentation, we propose a two-tier memory hierarchy: the recent past is stored in a short-term memory, where the storyboard temporal segments are produced and then transferred to a long-term memory. Experiments on two real-world egocentric video datasets show that contrastive learning objectives on top of storyboard frames result in semantically meaningful representations that outperform those produced by state-of-the-art unsupervised continual learning methods.
@@ 1 Introduction
Humans are capable of learning continuously from a stream of unlabeled and uncurated perceptual inputs, such as video data, without needing to iterate through multiple exposures or epochs. Since early infancy, humans have accumulated knowledge about the world through a continuous flow of raw visual observations. This capability contrasts sharply with the training paradigm of current methods in self-supervised learning (SSL) (Chen et al., 2020; Grill et al., 2020; Chen & He, 2021; Caron et al., 2020; Bardes et al., 2022; He et al., 2022; Assran et al., 2023; He et al., 2020). Despite making significant strides in learning from large unlabeled datasets, these approaches still predominantly rely on static and curated image datasets, such as ImageNet (Deng et al., 2009), and require multiple epochs of training for effective learning. This difference in paradigm raises a compelling question: how can we learn good visual representations in a streaming setting—learning from visual inputs in their original temporal order without cycling back?
Figure 1: Memory Storyboard framework for streaming self-supervised learning (SSL) from egocentric videos. Given a continuous stream of images from an egocentric video, Memory Storyboard effectively learns visual representations by clustering similar frames into temporal segments and updating their labels (text information for illustration purposes only) in the long-term memory buffer for replay. SSL involves contrastive learning at both the frame and temporal segment levels.
@@
Motivated by the differences in mechanisms between human learning and standard SSL, we aim to build learning algorithms that can efficiently learn visual representations and concepts from streaming video. One especially relevant mechanism in the human brain is event segmentation (Newtson et al., 1977; Zacks et al., 2001; Yates et al., 2022), where we spontaneously segment visual streams into hierarchically structured events and identify the event boundaries. Take your recent vacation trip as an example—you probably remember separate events and activities like exploring a city, dining at a local restaurant, or relaxing at the beach. The event segmentation mechanism helps us organize memories, recall specific moments, and summarize lengthened experiences (Zacks et al., 2006; Zacks & Swallow, 2007).
@@
Drawing inspiration from the way we organize our memory in the brain, we introduce Memory Storyboard, a novel approach for streaming self-supervised learning. Memory Storyboard features a temporal segmentation module, which groups video frames into semantically meaningful temporal segments, resembling the automatic event segmentation of human cognition. Through our temporal contrastive learning objective, these temporal segments effectively facilitate representation learning in streaming videos. To accommodate efficient temporal segmentation, we propose a two-tier hierarchical memory: temporal segmentation in the short-term memory is used to update the temporal class labels in the long-term memory, and a training batch consists of samples mixed from both memories. A high-level diagram of the algorithm is shown in Figure 1.
@@
We conduct experiments on the SAYCam (Sullivan et al., 2021) and KrishnaCam (Singh et al., 2016) datasets of real-world egocentric videos. Memory Storyboard outperforms state-of-the-art unsupervised continual learning methods on downstream image classification and object detection tasks and significantly reduces the gap between streaming learning and the less flexible IID learning that requires persistent storage of the entire prior video data. We also experiment with different buffer sizes and batch sizes and offer insights into the optimal training batch composition under different memory constraints.
@@
    We demonstrate that Memory Storyboard achieves state-of-the-art performance on downstream ImageNet (Deng et al., 2009) and iNaturalist (Van Horn et al., 2018) classification tasks when trained on real-world egocentric video datasets. Among all the streaming self-supervised learning methods we evaluated, Memory Storyboard is the only one that is competitive with or even outperforms IID training when trained on these datasets.
@@
    We study the effects of training factors including label merging, subsampling rate, average segment length, memory buffer size, and training batch composition. These studies provide insight for more efficient streaming learning from videos. In particular, we explore the optimal composition ratio of the training batch from short-term vs. long-term memory, under different memory constraints. Larger batches from long-term memory improve performance when we can afford a large memory bank, while smaller batches can help prevent overfitting when we have a small memory bank.
2 Streaming SSL from Egocentric Videos
In streaming self-supervised learning, the goal is to learn useful visual representations from a continuous stream of inputs $(x_{1},x_{2},\dots)$. Similar to continual learning, we impose a memory budget so that storing the entire video would violate the constraint. Different from standard continual learning, there is no explicit notion of task, and the data distribution shift follows directly from the scene transitions of a video. The learner needs to make changes to the model as it sees new inputs, and finishes learning as soon as it receives the last input of the stream. The streaming setting is similar to Online Continual Learning (Mai et al., 2021; Guo et al., 2022; Wei et al., 2023), but the focus here is primarily on streaming video frames instead of a fixed dataset of static images. We argue that streaming learning from sequential video frames enables better modeling of naturalistic scene transitions in real-world data streams because a stream of image collections often includes artificial class transitions.
Streaming Training Batches.
At each training step $t$, the model fetches a new batch of $b$ images $X_{t}=x_{tb:(t+1)b}$ from the video stream and updates its parameters upon receiving $X_{t}$. At the end of the video, we evaluate the final model checkpoint on various downstream tasks such as object classification and detection, which are fundamental tasks for visual scene understanding as they enable models to recognize and interpret the contents of complex environments.
Standard SSL Fails on Streaming Video.
Directly applying the SSL method sequentially on $X_{t}$ gives very poor performance (Purushwalkam et al., 2022; Ren et al., 2021). This is not only due to catastrophic forgetting (McCloskey & Cohen, 1989) caused by the non-stationary distribution of visual features in the stream, but also due to the high temporal correlation of images in the stream (illustrated in Figure 1). This temporal correlation breaks the IID assumption held by common optimization algorithms like SGD or Adam (Kingma & Ba, 2015). For contrastive learning algorithms like SimCLR (Chen et al., 2020), the similarity across different frames in the same training batch would violate the assumption that each image is different.
Memory Replay.
Similar to previous works (Hu et al., 2022; Yu et al., 2023; Purushwalkam et al., 2022), we use a replay buffer $M$ with finite size $|M|$ to mitigate these issues. The model can store some of the fetched images in the replay buffer, and use both samples from the replay buffer and the new frames to form a training batch of size $B$. By sampling from the replay buffer we de-correlate the frames in the training batch and at the same time reduce the distribution shift between training batches.
Benefits of Streaming SSL over Other Settings.
Compared to the traditional self-supervised learning setting, where all the frames are shuffled and uniformly sampled for each batch (we refer to this as "IID learning" in the text below), streaming SSL allows embodied agents to learn good visual representations from natural, uncurated video streams. It also involves less computation delay and less memory storage. For instance, a robot in a new environment can continuously adapt the visual representations from its own egocentric video feed without any human curation.
@@ 3 Related Work
In this section, we discuss the most relevant prior works. Please refer to Appendix B for additional related work.
@@
Unsupervised Continual Learning (UCL) (Rao et al., 2019; Smith et al., 2021; Madaan et al., 2022; Fini et al., 2022; Gomez-Villa et al., 2022; 2024; Cheng et al., 2023; Zhang et al., 2024) aims at learning a good representation through an unlabeled non-stationary data stream. Existing works in UCL often assume that the data stream is composed of a series of episodes and a stationary data distribution within each episode. This is not as naturalistic and human-like as our streaming setting, where the data distribution changes continuously through the data stream, and each image appears in the data stream only once. Meanwhile, we showed that existing UCL methods are also effective in our streaming video setting, and can be used together with the supervised contrastive objective.
@@
While a number of recent papers have studied streaming learning from images (Hayes et al., 2019; Hayes & Kanan, 2020; Hayes et al., 2020; Banerjee et al., 2021) or IID self-supervised learning from video frames (Venkataramanan et al., 2023; Wang et al., 2024), limited works have investigated the problem of streaming learning from a continuous video stream. Roady et al. (2020) introduces a benchmark for streaming classification and novelty detection from videos. Zhuang et al. (2022) benchmarks many self-supervised learning methods in real-time and life-long learning settings in streaming video, assuming infinite replay buffer size which is unrealistic. Most similar to our setup, Purushwalkam et al. (2022) studies the task of continuous representation learning with a SimSiam objective (Chen & He, 2021) and proposes using a minimum-redundancy replay buffer. Their work also belongs to the broader range of works that study replay buffer sampling strategies in continual learning (Aljundi et al., 2019; Wiewel & Yang, 2021; Tiwari et al., 2022; Hacohen & Tuytelaars, 2024). Our work extends these prior works by adopting a two-tier replay buffer and a temporal segmentation component. Also relevant to our work, Carreira et al. (2024) studies online learning from a continuous video stream using a pixel-to-pixel reconstruction loss for representation learning. Their findings on the effect of pre-training and different optimization schemes are orthogonal with the ones in our work. It is worth pointing out that their exploration mainly focuses on settings without data augmentation and replay, limiting the efficacy of their framework.
@@ 4 Memory Storyboard
We present Memory Storyboard, an effective method for streaming SSL from egocentric videos. Memory Storyboard includes a temporal segmentation module and a two-tier memory hierarchy. It combines a standard self-supervised contrastive loss with a temporal contrastive objective that leverages the temporal class labels produced by the temporal segmentation module. Figure 2 illustrates the details of our method. The overall data processing and training procedure is summarized in Algorithm 2.
@@
We describe our temporal segmentation algorithm as follows. Similar to Potapov et al. (2014), we are given a down-sampled video frame sequence of length $L$, with frames $x_{1},x_{2},\cdots,x_{L}$, and a feature extractor $f_{\theta}$. We aim to find change points $t_{1},t_{2},\cdots,t_{n-1}$ so that the video is divided into $n$ semantically-consistent segments $x_{1},x_{t_{1}}],[x_{t_{1}},x_{t_{2}}],\cdots,[x_{t_{n-1}},x_{L}]$. We also define $t_{0}=0$ and $t_{n}=L$. In this work, we determine the number of segments with $n=\frac{L}{T}$, where $T$ refers to the average segment length and is a hyper-parameter.
@@
The optimization objective of our segmentation algorithm is to maximize the average within-class similarity, such that each temporal segment captures a coherent scene, i.e.
$\max_{t_{1},t_{2},\cdots,t_{n-1}}\sum_{i=2}^{n}\frac{1}{t_{i}-t_{i-1}}\sum_{j=t_{i-1}}^{t_{i}}\sum_{k=j}^{t_{i}}sim(x_{j},x_{k}).$ (1)
@@
where $sim(x_{j},x_{k})$ denotes the cosine similarity between the embeddings $f_{\theta}(x_{j})$ and $f_{\theta}(x_{k})$. We compute the approximate solution to this optimization problem with a greedy approach, as detailed in Algorithm 1. We adopt this simple temporal segmentation approach in order to get good segmentation results in the beginning when the encoder network does not provide good representations. We leave it to future work to investigate different temporal segmentation strategies.
Figure 2: Details of our two-tier memory in Memory Storyboard. Long-term memory is updated with reservoir sampling and short-term memory with first-in-first-out (FIFO). Temporal segmentation is applied on the short-term memory, which then updates the labels of corresponding images in the long-term memory.
Two-tier Memory Hierarchy.
Inspired by the Complementary Learning Systems (CLS) theory (McClelland et al., 1995; O’Reilly et al., 2014) of the human brain, we propose a two-tier memory hierarchy to accommodate efficient temporal segmentation. Shown in Figure 2, the system includes a long-term memory $M_{long}$ updated with reservoir sampling (Vitter, 1985), and a short-term memory storyboard $M_{short}$ updated with a first-in-first-out (FIFO) strategy. We store the temporal index and the temporal class of each frame along with the image in the memory. The short-term memory size $|M_{short}|$ is much smaller than the long-term memory size $|M_{long}|$, allowing efficient temporal segmentation of the recent past. The change points produced by the temporal segmentation component on $M_{short}$ are then used to update the temporal class labels in $M_{long}$.
To increase the horizon of the memory storyboard, we subsample the frames coming from the current stream before adding it to $M_{short}$. The subsampling also reduces the temporal correlation between the frames in the training batch sampled from $M_{short}$. Compared to using a single replay buffer as memory, the two-tier memory hierarchy helps avoid overfitting the replay buffer and makes sure that the new frames are seen by the model.
Label Merging.
Same objects and scenes often repeat in egocentric video streams. To efficiently share visual concept labels, we introduce here a label merging mechanism. When a new temporal segment is added to $M_{\text{long}}$, we compute the cosine similarity between its average frame embedding and those of existing segments. If the maximum similarity exceeds a threshold $\delta$, the new segment inherits the class label of the most similar segment. This mechanism is activated only after the first $C$ segments, as early-stage embeddings tend to be uniformly high in similarity. Formally, let $v_{i}$ denote the average embedding of segment $i$ in $M_{\text{long}}$, and $v_{n}$ for the new segment $n$. Define $j=\arg\max_{i}\text{sim}(v_{i},v_{n})$ and let $c_{j}$ be the label of segment $j$. Then,
$c_{n}=\begin{cases}c_{j}&\text{if }\text{sim}(v_{j},v_{n})>\delta\text{ and }n>C\\ \text{new label}&\text{otherwise}.\end{cases}$
In practice, choosing a fixed threshold $\delta$ that generalizes across methods and datasets is challenging. To address this, rather than fixing $\delta$ manually, we define it dynamically based on a quantile threshold $\tau\in(0,1)$. Specifically, we set $\delta$ as the $\tau$-quantile of all off-diagonal values in the similarity matrix.
Temporal Contrastive Loss.
[…]
We refer to this as the temporal contrastive loss. It is conceptually similar to the temporal classification loss proposed in (Orhan et al., 2020). However, in the temporal classification loss, the size of the classification layer needs to be gradually expanded as more data is processed by the model and more temporal classes are formed. Hence, the temporal contrastive loss is more flexible and more suitable for the streaming SSL setting.
Overall Loss Function.
In addition to the temporal contrastive loss, we also incorporate a standard self-supervised loss $\mathcal{L}_{SSL}$. In particular, we experimented with the SimCLR loss (Chen et al., 2020; Sohn, 2016) and the SimSiam loss (Chen & He, 2021) because they were shown to work well in lifelong self-supervised learning in prior works (Zhuang et al., 2022; Purushwalkam et al., 2022). The overall loss function is a sum of the temporal contrastive loss and the self-supervised contrastive loss $\mathcal{L}=\mathcal{L}_{TCL}+\mathcal{L}_{SSL}$.
Warm-Start Training.
At the beginning of training, the model has only seen a very limited amount of data from the video stream. Even with a memory buffer, there is a likely high temporal correlation between the sampled frames which can cause instability in the training. To alleviate this problem, we warm-start the system by making no model updates on the first $M_{long}$ frames of the stream and just use them to fill the memory. The warm-start phase ensures that the model is trained on de-correlated samples from the buffer starting from the beginning.
@@
loss = TCL_loss(data) + SSL_loss(data)